The owner of a firm expects to make a profit of $100 for each of the next two year and to be able to sell the firm at the end of the second year for $800. The owner of the firm believes that the appropriate discount rate for the firm is 15%. Calculate the value of the firm!


1
Expert's answer
2021-02-03T02:48:57-0500

"\frac{100}{1+0.15}+\frac{100}{(1+0.15)^2}+\frac{800}{(1+0.15)^2}=767.49"